OpenStudy (anonymous):

How much should be deposited in an account paying 4.5% interest compounded monthly in order to have a balance of $11,000 after 3 years?

OpenStudy (anonymous):

Compound interest formula: FV=PV(1+i)^n. Where FV=future value, PV=present value, i=interest rate, n=number of periods compounded. Remember you can manipulate this formula to find which value you are looking for. In your question we're looking for PV. Dividing both sides of the equation above by (1+i)^n we get PV=FV/((1+i)^n). After this, all we have to do is plug in what we have:\[PV = 11000/((1+0.045)^{12*3})\] Note 12*3 is how many months total. Solving that will get you the answer.

OpenStudy (anonymous):

And you can see how much money you can make on just interest! ...credit card companies use that to their advantage.

OpenStudy (anonymous):

blahhh im still stuck.

OpenStudy (anonymous):

What's giving you troubles

OpenStudy (anonymous):

that last equation >.<

OpenStudy (anonymous):

The number or how I got to that equation?

OpenStudy (anonymous):

how to figure out the answer with the equation. >.^

OpenStudy (anonymous):

Try doing it in parts. Start with the bottom. 1+.045 is obviously 1.045 (remember 4.5%=0.045). Then do 1.045^36. Write that number down. I got 4.8774. Then do 11000/4.8774

OpenStudy (anonymous):

 Is that easier to understand?

OpenStudy (anonymous):

yea a lil